A perfectly price discriminating monopoly has 2 consumers: consumer 1 with demand curve q1=10-p and consumer 2 with demand curve q2=8-p. The monopoly's cost function is c(Q)=Q2. The quantity consumer 2 will consume is:
1
Expert's answer
2020-10-16T10:43:49-0400

The quantity consumer 2 will consume is:

Q = q1 + q2 = 18 - p,

p = MC,

p = 18 - Q,

MC = C'(Q) = 2Q,

18 - Q = 2Q,

Q = 6 units.

8 - q2 = 2Q,

q2 = 2.67 units.